Circular No. 6829

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998S IN NGC 3877 W.-d. Li, University of California at Berkeley, forwards a report from

C. Li, Beijing Astronomical Observatory (BAO), that Zhou Wan has found an apparent supernova (unfiltered mag 15.2) on CCD images obtained with the 0.60-m BAO telescope on Mar. 3 UT as part of the BAO Supernova Survey (cf. IAUC 6612). A confirming CCD observation made with the Katzman Automatic Imaging Telescope (KAIT) of Lick Observatory by the Lick Supernova Survey (cf. IAUC 6627) indicates that the object has brightened to unfiltered mag ∼ 13.5 on Mar. 4.3 UT. SN 1998S is located at α = 11^{h}46^{m}06^{s}, δ = +47^{o}29′.0 (equinox 2000.0; based on the KAIT image), which is 16′′ west and 46′′ south of the nucleus of NGC 3877. As the object is embedded in the galaxy, the BAO magnitude was obtained from a subtracted image. SN 1998S is not present in images of the same field taken prior to Mar. 3.

A. V. Filippenko and E. C. Moran, University of California at Berkeley, obtained a very-high-resolution (0.015 nm) échelle spectrogram of SN 1998S on Mar. 4 with the Keck-1 telescope. The probable presence of broad Hα emission superposed on a featureless continuum suggests that this is a type-II supernova.

PSR J1024–0719 W. Becker and J. Trümper, Max-Planck-Institut für Extraterrestrische

Physik, report: “We have detected the isolated galactic millisecond pulsar PSR J1024–0719 in the energy range 0.1–2.4 keV with ROSAT. The source was observed with the High Resolution Imager (HRI) from 1997 Nov. 17 to Dec. 7 for a total exposure time of 80 635 s. The net countrate (including background, vignetting and dead-time corrections) is (3.3 ± 0.3) × 10^{−4} counts/s. Assuming a Crab-like spectrum and a column absorption of N_{H} = 2 × 10^{20} cm^{−2}, deduced from the radio dispersion measure (cf. Bailes et al. 1997, Ap.J. 481, 386), we find an energy flux of 2 × 10^{−14} erg s^{−1} cm^{−2} over 0.1–2.4 keV. The low number of recorded source counts precluded any modulation testing at the 5.15-ms radio period.”

COMET 69P/TAYLOR Visual m_{1} estimates: Feb. 14.83 UT, 11.3 (A. Kammerer, Ettlingen,

Germany, 0.20-m reflector); 27.05, 11.8 (J. Bortle, Stormville, NY, 0.41-m reflector); Mar. 1.78, 11.8 (V. Znojil, Brno, Czech Rep., 25×100 binoculars).

1998 March 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Daniel W. E. Green